OpenStudy (anonymous):

3x^2 + 8x=5x^2-2x solve for x ... please help, these exponents are throwing me off!!

OpenStudy (jkristia):

move it all to one side of the equal sign and set it = 0, then you have 3x^2 + 8x - 5x^2 - 2x = 0 -2x^2 + 10x = 0 now solve for x

OpenStudy (jkristia):

it is "+2x", not "-2x"

OpenStudy (anonymous):

Thank you! But I'm still having trouble getting x.. how do i get rid of the exponent?

OpenStudy (jkristia):

-2x^2 + 10, now try factor -2x out, so -2x * (something in here) = 0 what do need to multiply -2x by to get -2x^2 + 10x ?

OpenStudy (anonymous):

-2x..right?

OpenStudy (anonymous):

wait no...

OpenStudy (jkristia):

-2x * x = -2x^2 -2x * -5 = +10x so -2x * x + -2x * - 5 => -2x * (x - 5) = 0